Grade 11Mechanics

A stone is thrown from a bridge at an angle of 30° down with the horizontal with a velocity of 25m/s. If the stone strikes the water after 2.5s, calculate the height of the bridge from the water surface.

Durwa Thakur
8 Years agoGrade 11

2 Answers

Khimraj
ApprovedApproved Tutor Answer8 Years ago
vertical velocity = 25sin30  = 25/2
t = 5/2
then applying second equation of motion
h = ut + ½gt2
h = (25/2)*(5/2) + ½*10*(5/2)2
h = 125/4 + 125/4 = 125/2 = 62.5 m
Hope it clears. If you like answer then please approve it.
Yuvraj
6 Years ago
Correct answer is 61.9 as we see the line passing through top of bridge and the stone the velocity of stone at the point on line is +25/2 m/s and the acceleration is +g (as we take downward direction as positive ) so by distance formula s=ut+1/2at² we can find h